Eight students attended a birthday party: Three * (3) of them confirmed Couid d - 10 positive. The contact tracer bad conducted rondom. test for only two students. Find the probabi lity distribution for random variable & which represents the number of Covid-1a positive patients. Now, think of three stedements of the problem out of the topic "Covid-19?"
1
Expert's answer
2021-10-19T03:01:22-0400

This scenario can be described by a binomial distribution because the experiment results in an outcome that can be classified as either a success or failure. For this case, the test results in either a student being Covid-19 positive(success) or Covid-19 negative(failure).

Let "X" be a random variable representing the number of Covid-19 patients, "n" be the number of trials, "x" be the number of successes out of "n" trials and "p" be the probability of success.

Therefore, out of "n=8" students, the number of successes are "x=3" and the probability of success "p=(x\/n)=3\/8". The probability of failure is given as "q=1-p=1-3\/8=5\/8" .

The probability distribution of the random variable "X" is given by the formula,

"p(X=x)=\binom{8}{x}p^x(1-p)^{n-x},\space x=0,1,2,3,4,5,6,7,8"

"0,\space elsewhere"

If a random test was done on 2 students then, probability that they were both Covid-19 positive is given as,

"p(X=2)=\binom{8}{2}(3/8)^{2}*(5/8)^6=0.2346933"

The probability that both are Covid-19 negative is "1-0.2346933=0.7653067".